OpenStudy (anonymous):

HELP ME PLEASE!! 3x=6 = 7x-1 ----- ----- 8 6

OpenStudy (anonymous):

Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation.

OpenStudy (anonymous):

3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (anonymous):

Is the original equation "3x=6=7x-1"?

OpenStudy (anonymous):

the 3x=6 is over the 8 and the 7x-1 is over six but yes

OpenStudy (anonymous):

\[3x+6\over 8\]is equal to \[7x-1\over 6\]

OpenStudy (texaschic101):

That is a proportion...but I can't figure out why she set it up like that. I would do this.. (3x + 6) / 8 = (7x - 1) / 6 cross multiply 6(3x + 6) = 8(7x - 1) I don't know where she is coming up with the 3 and the 4 that are before the parenthesis

OpenStudy (anonymous):

me either

OpenStudy (anonymous):

but i need to know were they went wrong 3(3x+6)=4(7x−1) 9x+18=28x−1 9x−9x+18=28x−9x−1 18=19x−1 18+1=19x−1+1 19=19x 1=x

OpenStudy (texaschic101):

the answer can be checked by subbin what she got for x, which is 1, back into the equation and see if it is correct.

OpenStudy (anonymous):

all the question marks are subtractions and negitives

OpenStudy (anonymous):

why is it saying i was typing

OpenStudy (anonymous):

and at the very top its 3x+6 ----- 8

OpenStudy (anonymous):

hello?

OpenStudy (texaschic101):

wait....I think I see (3x + 6) / 8 = (7x - 1) / 6 cross multiply 6(3x + 6) = 8(7x - 1) -- she reduced the6 and the 8 by dividing by 2 3(3x + 6) = 4(7x - 1) ==== < first error..when she distributed, she didn't multiply the -1 by 4 correctly..it should come out to what I got below 9x + 18 = 28x - 4 -- subtract 9x from both sides 18 = 28x - 9x - 4 -- add 4 to both sides 18 + 4 = 28x - 9x -- combine like terms 22 = 19x 22/19 = x

OpenStudy (texaschic101):

that is the correct answer. She just distributed wrong

OpenStudy (anonymous):

thank you